what is the surface area of a cylinder with a radius of 3 and a height of 6?
Answer by lenny460(1073)

The radius = 3
The height = 6
The Surface Area of a Cylinder:
2(pi r^2) + (2 pi r)h o
or simply:
2pir^2 + 2pirh

Let's find out the area with the radius of 3
2pir^2 = 2 * 3.14 * 3 * 3
3 * 3 = 9
2 * 3.14 * 9 = 56.52

Now let's find out the area with the height of 6
2pirh = 2 * 3.14 * 3 * 6
3 * 6 = 18
2 * 3.14 * 18 = 113.04

We now add the two:
113.04 + 56.52 = 169.56

So The Answer could be:
169.56 cm^2 or
169.56 in^2
You did not indicate the measurement unit.


obuong@aol.com